A point charge with q1

2.5 C is located at x  3.0 m,

y  0, and a second point charge

with q2  4.0 C is at x  1.0 m,

y  2.0 m. (a) What is the force

exerted by q1 on q2? (b) What is the

force exerted by q2 on q1?
Answer #1

a)Kq1q2/r^2=9*10^9*2.5*4/8=1.125*10^10 N

since r=((3-1)^2+(0-2)^2)^1/2

b)q2 will exert same force on q1 but in opposite direction
Answer #2

the charge q1 = 2.5 C is at (3,0) m

the charge q2 = 4.0 C is at (1,2) m

the force between q1 and q2 is

F = k x (q1q2/r^2)

where k = (1/4pi x e_o) = 9 x 10^9 Nm^2/C^2 and r^2 = [(3 -1)^2 + (0 - 2)^2] m^2